# Optimal rates of approximation by shallow ReLUk neural networks and applications to nonparametric regression

### 21 Jun 2023 v2

@@ 2 Approximation rates for shallow neural networks
(2) Given $h\in\mathcal{H}^{\alpha}$, for any $u=(u_{1},\dots,u_{d+1})^{\intercal}\in\Omega$, we define $\widetilde{h}(u):=u_{d+1}^{k}h(u_{d+1}^{-1}u^{\prime})$, where $u^{\prime}=(u_{1},\dots,u_{d})^{\intercal}$. It is easy to check that $S_{k}\widetilde{h}=h$. Note that $h$ is completely determined by the function values of $\widetilde{h}$ on $\Omega$. Observe that the smoothness of $\widetilde{h}$ on $\Omega$ can be controlled by the smoothness of $h$. We can extend $\widetilde{h}$ to $\mathbb{R}^{d+1}$ so that $\|\widetilde{h}\|_{C^{r,\beta}(\mathbb{R}^{d+1})}\leq C_{0}$ for some constant $C_{0}$ independent of $h$, by using (refined version of) Whitney’s extension theorem [Fefferman, 2006, 2009; Fefferman and Israel, 2020]. It remains to show that $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}\lesssim t^{\alpha}$.
@@
$\omega_{2s^{*}}(\widetilde{h},t)_{p}$ $\lesssim\sup_{0<\theta\leq t}\sup_{u\in\mathbb{S}^{d}}\left|\sum_{j=0}^{2s^{*}}(-1)^{j}\binom{2s^{*}}{j}\int_{\mathbb{S}^{\perp}_{u}}\widetilde{h}(u\cos j\theta+v\sin j\theta)d\tau_{d-1}(v)\right|$
@@
Next, we estimate $\sup_{0<\theta\leq t}|H(u,v,\theta)|$ for small $t>0$ and fixed $u,v$. One can check that the function $f(\cdot):=\widetilde{h}(u\cos(\cdot)+v\sin(\cdot))$ is in $C^{r,\beta}([0,t_{0}])$ for small $t_{0}>0$, and $\|f\|_{C^{r,\beta}([0,t_{0}])}\lesssim\|\widetilde{h}\|_{C^{r,\beta}(\mathbb{R}^{d+1})}$. Let $\widetilde{\Delta}_{\theta}f(\cdot):=f(\cdot+\theta)-f(\cdot)$ be the difference operator and $\widetilde{\Delta}_{\theta}^{n+1}:=\widetilde{\Delta}_{\theta}\widetilde{\Delta}_{\theta}^{n}$ for $n\in\mathbb{N}$. The binomial theorem shows $H(u,v,\theta)=\widetilde{\Delta}_{\theta}^{2s^{*}}f(0)$. Then, classical theory of moduli of smoothness [DeVore and Lorentz, 1993, Chapter 2.6–2.9] implies
@@
By Proposition 3.3, for any $h\in\mathcal{H}^{\alpha}$, there exists $\widetilde{h}\in C(\mathbb{S}^{d})$ such that $S_{k}\widetilde{h}=h$, $\|\widetilde{h}\|_{L^{\infty}(\mathbb{S}^{d})}\leq C$ and $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}\leq Ct^{\alpha}$, where $s^{*}\in\mathbb{N}$ is the smallest integer such that $\alpha\leq 2s^{*}$. We choose $\widetilde{h}$ to be odd (even) if $k$ is even (odd). Using $\omega_{s}(\widetilde{h},t)_{2}\leq 2^{s-2s^{*}+2}\omega_{2s^{*}}(\widetilde{h},t)_{2}$ for $s>2s^{*}$ [Dai and Xu, 2013, Proposition 10.1.2] and the Marchaud inequality [Ditzian, 2006, Eq.(9.6)]
@@
which converges in $L^{2}(\mathbb{S}^{d})$. This implies that $g\in\mathcal{F}_{\sigma_{k}}(M)$ if $\mathcal{P}_{n}g=0$ for any $n\in\mathbb{N}_{0}$ satisfying $\widehat{\sigma_{k}}(n)=0$ and
@@
where $\eta$ is a $C^{\infty}$-function on $0,\infty)$ such that $\eta(t)=1$ for $0\leq t\leq 1$ and $\eta(t)=0$ for $t\geq 2$. Since $\eta$ is supported on $[0,2]$, the summation can be terminated at n=2m1n=2m-1, so that $g_{m}$ is a polynomial of degree at most 2m12m-1. Since $\widetilde{h}$ is odd (even) if $k$ is even (odd), $\mathcal{P}_{n}g_{m}=\eta(n/m)\mathcal{P}_{n}\widetilde{h}=0$ for any $n\equiv k\bmod 2$. Furthermore, Dai and Xu [2013, Theorem 10.3.2] shows that
$K_{s}(\widetilde{h},m^{-1})_{p}\asymp\|\widetilde{h}-g_{m}\|_{L^{p}(\mathbb{S}^{d})}+m^{-s}\|(-\Delta)^{s/2}g_{m}\|_{L^{p}(\mathbb{S}^{d})}.$
By the equivalence (3.4) and $\omega_{2s^{*}}(\widetilde{h},m^{-1})_{\infty}\lesssim m^{-\alpha}$, we can bound the approximation error as
@@
where we choose $s=(d+2k+1)/2$ in the last inequality. There are three cases:
@@
$\leq\lim_{m\to\infty}\sum_{\widehat{\sigma_{k}}(n)\neq 0}\widehat{\sigma_{k}}(n)^{-2}\|\mathcal{P}_{n}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})}\lesssim 1.$
This shows that $\widetilde{h}\subseteq\mathcal{G}_{\sigma_{k}}(M)$ for some constant $M$. Hence, $h=S_{k}\widetilde{h}\in\mathcal{F}_{\sigma_{k}}(M)$ by Proposition 3.3.
@@
Since we are only able to estimate the smoothness $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}$ for even integer $2s^{*}$, we have an extra logarithmic factor for the bound $\omega_{s}(\widetilde{h},t)_{2}\lesssim t^{\alpha}\sqrt{\log(1/t)}$ in (3.5) when $s=\alpha\neq 2s^{*}$, due to the Marchaud inequality. Consequently, we can only obtain exponential convergence rate when $\alpha=(d+2k+1)/2$ is not an even integer. We conjecture the bound $\omega_{\alpha}(\widetilde{h},t)_{2}\lesssim t^{\alpha}$ holds for all $\alpha>0$. If this is the case, then the proof of Theorem 2.1 implies $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma_{k}}(M)$ for some constant $M$ when $\alpha\geq(d+2k+1)/2$.
@@ 4 Nonparametric regression
In this section, we apply our approximation results to nonparametric regression using neural networks. For simplicity, we will only consider ReLU activation function ($k=1$), which is the most popular activation in deep learning. To use the uniform approximation bound in Corollary 2.4, we also require that the dimension $d\geq 4$.
We consider the classical problem of learning a $d$-variate function $h\in\mathcal{H}$ from its noisy samples, where we will assume $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$ or $\mathcal{H}=\mathcal{F}_{\sigma}(1)$. Note that, due to Theorem 2.1, the results for $\mathcal{F}_{\sigma}(1)$ can be applied to $\mathcal{H}^{\alpha}$ with $\alpha>(d+3)/2$ by scaling the variation norm. Suppose we have a data set of $n\geq 2$ samples $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}\subseteq\mathbb{B}^{d}\times\mathbb{R}$ which are independently and identically generated from the regression model
@@
where $\mathcal{F}_{n}$ is a function class parameterized by neural networks. For simplicity, we assume here and in the sequel that the minimum above indeed exists. When this is not the case, our results also hold for any estimator which minimizes the empirical risk up to a small error term. The performance of the estimation is measured by the expected risk
@@
It is equivalent to evaluate the estimator by the excess risk
@@ 4.1 Shallow neural networks
The rate of convergence of neural network regression estimates has been analyzed by many papers [McCaffrey and Gallant, 1994; Kohler and Krzyżak, 2005; Bauer and Kohler, 2019; Schmidt-Hieber, 2020; Nakada and Imaizumi, 2020; Kohler and Langer, 2021]. It is well-known that the optimal minimax rate of convergence for learning a regression function $h\in\mathcal{H}^{\alpha}$ is $n^{-2\alpha/(d+2\alpha)}$ [Stone, 1982]. This optimal rate has been established (up to logarithmic factors) for two-hidden-layers neural networks with certain squashing activation functions [Kohler and Krzyżak, 2005] and for deep ReLU neural networks [Schmidt-Hieber, 2020; Kohler and Langer, 2021]. For shallow networks, McCaffrey and Gallant [1994] proved a rate of $n^{-2\alpha/(2\alpha+d+5)+\epsilon}$ for a certain cosine squasher activation function. However, to the best of our knowledge, it is unknown whether shallow neural networks can achieve the optimal rate. In this section, we provide an affirmative answer to this question by proving that shallow ReLU neural networks can achieve the optimal rate for $\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$.
@@
For shallow neural network model $\mathcal{F}_{n}=\mathcal{F}_{\sigma}(N_{n},M_{n})$, Lemma 2.3 and Corollary 2.4 provide bounds for the approximation errors. The covering number of $\mathcal{F}_{n}$ can be estimated by using the Lipschitz continuity of the activation function, since the weights in the network are bounded. Choosing $N_{n},M_{n}$ appropriately to balance the approximation and generalization errors, we can derive convergence rates for the ERM.
@@
Let $d\geq 4$, $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\mathcal{F}_{\sigma}(N_{n},M_{n})$ and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$.
@@
    $N_{n}\asymp n^{\frac{d}{d+2\alpha}},\quad M_{n}\asymp n^{\frac{d+3-2\alpha}{2d+4\alpha}},$
@@
    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}(\log n)^{3}.$
@@
    $N_{n}\asymp n^{\frac{d}{2d+3}},\quad M_{n}=M,$
    for some sufficiently large constant $M\geq 1$, then
    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{d+3}{2d+3}}(\log n)^{3}.$
@@
To apply the bound in Lemma 4.1, we need to estimate the covering number $\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})$. It is easy to check that $\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\leq\mathcal{N}(\epsilon,\mathcal{F}_{n},\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})$. Hence, it is enough to provide a bound for $\mathcal{N}(\epsilon,\mathcal{F}_{n},\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})$.
Recall that any $f\in\mathcal{F}_{n}$ can be represented in the form $f(x)=\sum_{i=1}^{N_{n}}a_{i}\sigma((x^{\intercal},1)v_{i})$ with $v_{i}\in\mathbb{S}^{d}$ and $\sum_{i=1}^{N_{n}}|a_{i}|\leq M_{n}$. It is easy to see that there exists $\epsilon/(2\sqrt{2}M_{n})$-cover $V_{\epsilon}$ of $\mathbb{S}^{d}$ such that its cardinality $|V_{\epsilon}|\lesssim(1+M_{n}/\epsilon)^{d}$. Similarly, there exists $\epsilon/(2\sqrt{2})$-cover $A_{\epsilon}$ of the set $A=\{a=(a_{1},\dots,a_{N_{n}}):\|a\|_{1}\leq M_{n}\}$ in the metric $\|\cdot\|_{1}$ such that $|A_{\epsilon}|\lesssim(1+M_{n}/\epsilon)^{N_{n}}$. Hence, for any $v_{i}\in\mathbb{S}^{d}$ and $a=(a_{1},\dots,a_{N_{n}})\in A$, there exists $\widetilde{v}_{i}\in V_{\epsilon}$ and $\widetilde{a}=(\widetilde{a}_{1},\dots,\widetilde{a}_{N_{n}})\in A_{\epsilon}$ satisfying $\|v_{i}-\widetilde{v}_{i}\|_{2}\leq\epsilon/(2\sqrt{2}M_{n})$ and $\|a-\widetilde{a}\|_{1}\leq\epsilon/(2\sqrt{2})$. Therefore, letting $\widetilde{f}(x)=\sum_{i=1}^{N_{n}}\widetilde{a}_{i}\sigma((x^{\intercal},1)\widetilde{v}_{i})$, we have
$|f(x)-\widetilde{f}(x)|$ $\leq\|a-\widetilde{a}\|_{1}\sup_{1\leq i\leq N_{n}}|\sigma((x^{\intercal},1)v_{i})|+\|\widetilde{a}\|_{1}\sup_{1\leq i\leq N_{n}}|\sigma((x^{\intercal},1)v_{i})-\sigma((x^{\intercal},1)\widetilde{v}_{i})|$
$\leq\epsilon/2+\sqrt{2}M_{n}\sup_{1\leq i\leq N_{n}}\|v_{i}-\widetilde{v}_{i}\|_{2}\leq\epsilon,$
where we use the Lipschitz continuity of ReLU in the second inequality. This shows that
$\log\mathcal{N}(\epsilon,\mathcal{F}_{n},\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})\lesssim N_{n}\log|V_{\epsilon}|+\log|A_{\epsilon}|\lesssim N_{n}\log(1+M_{n}/\epsilon).$
@@
$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{(\log n)^{2}N_{n}\log(nM_{n}B_{n})}{n}+N_{n}^{-\frac{2\alpha}{d}}\lor M_{n}^{-\frac{4\alpha}{d+3-2\alpha}}.$
By choosing $N_{n}\asymp n^{d/(d+2\alpha)}$ and $M_{n}\asymp N_{n}^{(d+3-2\alpha)/(2d)}$, we get $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-2\alpha/(d+2\alpha)}(\log n)^{3}$. Similarly, by Lemma 4.1 and Lemma 2.3, if $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, there exists $M_{n}=M$ such that
$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{(\log n)^{2}N_{n}\log(nM_{n}B_{n})}{n}+N_{n}^{-\frac{d+3}{d}}.$
We choose $N_{n}\asymp n^{d/(2d+3)}$, then $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-(d+3)/(2d+3)}(\log n)^{3}$. ∎
@@
The Gaussian noise assumption on the model (4.1) can be weaken for Lemma 4.1 and hence for Theorem 4.2. We refer the reader to Kohler and Langer [2021, Appendix B, Lemma 18] for more details. Theorem 4.2 can be easily generalized to shallow ReLUk neural networks for $k\geq 1$ by using the same proof technique, if we assume that $\mu$ has a bounded density (so that we can use Corollary 2.4 to bound the approximation error in the $L^{2}(\mu)$ norm). For example, one can show that, if $h\in\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$, then we can choose $\mathcal{F}_{n}=\mathcal{F}_{\sigma_{k}}(N_{n},M_{n})$ with $N_{n}\asymp n^{\frac{d}{d+2\alpha}}$ and $M_{n}\asymp n^{\frac{d+2k+1-2\alpha}{2d+4\alpha}}$, such that $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}(\log n)^{3}$.
Theorem 4.2 shows that least square minimization using shallow ReLU neural networks can achieve the optimal rate $n^{-\frac{2\alpha}{d+2\alpha}}$ for learning functions in $\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$. For the function class $\mathcal{F}_{\sigma}(1)$, the rate $n^{-\frac{d+3}{2d+3}}$ is also minimax optimal as proven by Parhi and Nowak [2023, Lemma 25] (they studied a slightly different function class, but their result also holds for $\mathcal{F}_{\sigma}(1)$). Specifically, Siegel and Xu [2022, Theorem 4 and Theorem 8] give a sharp estimate for the metric entropy
@@
Combining this estimate with the classical result of Yang and Barron (see [Yang and Barron, 1999, Proposition 1] and [Wainwright, 2019, Chapter 15]), we get
@@
where the infimum is over all functions of the samples $\mathcal{D}_{n}$, which are generated from the model (4.1).

### 09 Jan 2024 v3

@@ 2 Approximation rates for shallow neural networks
(2) Given $h\in\mathcal{H}^{\alpha}$, for any $u=(u_{1},\dots,u_{d+1})^{\intercal}\in\Omega$, we define $\widetilde{h}(u):=u_{d+1}^{k}h(u_{d+1}^{-1}u^{\prime})$, where $u^{\prime}=(u_{1},\dots,u_{d})^{\intercal}$. It is easy to check that $S_{k}\widetilde{h}=h$. Note that $h$ is completely determined by the function values of $\widetilde{h}$ on $\Omega$. Observe that the smoothness of $\widetilde{h}$ on $\Omega$ can be controlled by the smoothness of $h$. We can extend $\widetilde{h}$ to $\mathbb{R}^{d+1}$ so that $\|\widetilde{h}\|_{C^{r,\beta}(\mathbb{R}^{d+1})}\leq C_{0}$ for some constant $C_{0}$ independent of $h$, by using (refined version of) Whitney’s extension theorem [17, 18, 19]. It remains to show that $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}\lesssim t^{\alpha}$.
@@
$\omega_{2s^{*}}(\widetilde{h},t)_{\infty}$ $\lesssim\sup_{0<\theta\leq t}\sup_{u\in\mathbb{S}^{d}}\left|\sum_{j=0}^{2s^{*}}(-1)^{j}\binom{2s^{*}}{j}\int_{\mathbb{S}^{\perp}_{u}}\widetilde{h}(u\cos j\theta+v\sin j\theta)d\tau_{d-1}(v)\right|$
@@
Next, we estimate $\sup_{0<\theta\leq t}|H(u,v,\theta)|$ for small $t>0$ and fixed $u,v$. One can check that the function $f(\cdot):=\widetilde{h}(u\cos(\cdot)+v\sin(\cdot))$ is in $C^{r,\beta}([0,t_{0}])$ for small $t_{0}>0$, and $\|f\|_{C^{r,\beta}([0,t_{0}])}\lesssim\|\widetilde{h}\|_{C^{r,\beta}(\mathbb{R}^{d+1})}$. Let $\widetilde{\Delta}_{\theta}f(\cdot):=f(\cdot+\theta)-f(\cdot)$ be the difference operator and $\widetilde{\Delta}_{\theta}^{n+1}:=\widetilde{\Delta}_{\theta}\widetilde{\Delta}_{\theta}^{n}$ for $n\in\mathbb{N}$. The binomial theorem shows $H(u,v,\theta)=\widetilde{\Delta}_{\theta}^{2s^{*}}f(0)$. Then, the classical theory of moduli of smoothness [14, Chapter 2.6–2.9] implies
@@
By Proposition 3.3, for any $h\in\mathcal{H}^{\alpha}$, there exists $\widetilde{h}\in C(\mathbb{S}^{d})$ such that $S_{k}\widetilde{h}=h$, $\|\widetilde{h}\|_{L^{\infty}(\mathbb{S}^{d})}\leq C$ and $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}\leq Ct^{\alpha}$, where $s^{*}\in\mathbb{N}$ is the smallest integer such that $\alpha\leq 2s^{*}$. We choose $\widetilde{h}$ to be odd (even) if $k$ is even (odd). Using $\omega_{s}(\widetilde{h},t)_{2}\leq 2^{s-2s^{*}+2}\omega_{2s^{*}}(\widetilde{h},t)_{2}$ for $s>2s^{*}$ [12, Proposition 10.1.2] and the Marchaud inequality [15, Eq.(9.6)]
@@
which converges in $L^{2}(\mathbb{S}^{d})$. This implies that $g\in\mathcal{G}_{\sigma_{k}}(M)$ if $g$ is continuous, $\mathcal{P}_{n}g=0$ for any $n\in\mathbb{N}_{0}$ satisfying $\widehat{\sigma_{k}}(n)=0$ and
@@
where $\eta$ is a $C^{\infty}$-function on $0,\infty)$ such that $\eta(t)=1$ for $0\leq t\leq 1$ and $\eta(t)=0$ for $t\geq 2$. Since $\eta$ is supported on $[0,2]$, the summation can be terminated at n=2m1n=2m-1, so that $g_{m}$ is a polynomial of degree at most 2m12m-1. Since $\widetilde{h}$ is odd (even) if $k$ is even (odd), $\mathcal{P}_{n}g_{m}=\eta(n/m)\mathcal{P}_{n}\widetilde{h}=0$ for any $n\equiv k\bmod 2$. Furthermore, [12, Theorem 10.3.2] shows that
$K_{s}(\widetilde{h},m^{-1})_{p}\asymp\|\widetilde{h}-g_{m}\|_{L^{p}(\mathbb{S}^{d})}+m^{-s}\|(-\Delta)^{s/2}g_{m}\|_{L^{p}(\mathbb{S}^{d})}.$ (3.6)
By the equivalence (3.4) and $\omega_{2s^{*}}(\widetilde{h},m^{-1})_{\infty}\lesssim m^{-\alpha}$, the equivalence (3.6) for $p=\infty$ implies that we can bound the approximation error as
@@
where we choose $s=(d+2k+1)/2$ in the last inequality. We continue the proof in three different cases.
@@
$\leq\lim_{m\to\infty}\sum_{\widehat{\sigma_{k}}(n)\neq 0}\widehat{\sigma_{k}}(n)^{-2}\|\mathcal{P}_{n}g_{m}\|^{2}_{L^{2}(\mathbb{S}^{d})}$
$\lesssim 1.$
This shows that $\widetilde{h}\subseteq\mathcal{G}_{\sigma_{k}}(M)$ for some constant $M$. Hence, $h=S_{k}\widetilde{h}\in\mathcal{F}_{\sigma_{k}}(M)$ by Proposition $3.$
@@
Since we are only able to estimate the smoothness $\omega_{2s^{*}}(\widetilde{h},t)_{\infty}$ for even integer $2s^{*}$, we have an extra logarithmic factor for the bound $\omega_{s}(\widetilde{h},t)_{2}\lesssim t^{\alpha}\sqrt{\log(1/t)}$ in (3.5) when $s=\alpha\neq 2s^{*}$, due to the Marchaud inequality. Consequently, we can only obtain exponential convergence rate when $\alpha=(d+2k+1)/2$ is not an even integer. We conjecture the bound $\omega_{s}(\widetilde{h},t)_{2}\lesssim t^{\alpha}$ holds for all $s\geq\alpha$. If this is the case, then the proof of Theorem 2.1 implies $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma_{k}}(M)$ for some constant $M$ when $\alpha\geq(d+2k+1)/2$.
@@ 4 Nonparametric regression
In this section, we apply our approximation results to nonparametric regression using neural networks. For simplicity, we will only consider ReLU activation function ($k=1$), which is the most popular activation in deep learning.
We study the classical problem of learning a $d$-variate function $h\in\mathcal{H}$ from its noisy samples, where we will assume $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$ or $\mathcal{H}=\mathcal{F}_{\sigma}(1)$. Note that, due to Theorem 2.1, the results for $\mathcal{F}_{\sigma}(1)$ can be applied to $\mathcal{H}^{\alpha}$ with $\alpha>(d+3)/2$ by scaling the variation norm. Suppose we have a data set of $n\geq 2$ samples $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}\subseteq\mathbb{B}^{d}\times\mathbb{R}$ which are independently and identically generated from the regression model
@@
where $\mathcal{F}_{n}$ is a function class parameterized by neural networks. For simplicity, we assume here and in the sequel that the minimum above indeed exists. The performance of the estimation is measured by the expected risk
@@
It is equivalent to evaluating the estimator by the excess risk
@@ 4.1 Shallow neural networks
The rate of convergence of neural network regression estimates has been analyzed by many papers [38, 26, 9, 52, 41, 27]. It is well-known that the optimal minimax rate of convergence for learning a regression function $h\in\mathcal{H}^{\alpha}$ is $n^{-2\alpha/(d+2\alpha)}$ [59]. This optimal rate has been established (up to logarithmic factors) for two-hidden-layers neural networks with certain squashing activation functions [26] and for deep ReLU neural networks [52, 27]. For shallow networks, [38] proved a rate of $n^{-2\alpha/(2\alpha+d+5)+\epsilon}$ with $\epsilon>0$ for a certain cosine squasher activation function. However, to the best of our knowledge, it is unknown whether shallow neural networks can achieve the optimal rate. In this section, we provide an affirmative answer to this question by proving that shallow ReLU neural networks can achieve the optimal rate for $\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$.
@@
For shallow neural network model $\mathcal{F}_{n}=\mathcal{F}_{\sigma}(N_{n},M_{n})$, Lemma 2.3 and Corollary 2.4 provide bounds for the approximation errors. The covering number of the function class $\mathcal{T}_{B_{n}}\mathcal{F}_{n}$ can be estimated by using the pseudo-dimension of $\mathcal{T}_{B_{n}}\mathcal{F}_{n}$ [22]. Choosing $N_{n},M_{n}$ appropriately to balance the approximation and generalization errors, we can derive convergence rates for the ERM.
@@
Let $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\mathcal{F}_{\sigma}(N_{n},M_{n})$ and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$.
@@
    $N_{n}\asymp n^{\frac{d}{d+2\alpha}},\quad M_{n}\gtrsim n^{\frac{d+3-2\alpha}{2d+4\alpha}},$
@@
    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}(\log n)^{4}.$
@@
    $N_{n}\asymp n^{\frac{d}{2d+3}},\quad M_{n}\geq 1,$
    then
    $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{d+3}{2d+3}}(\log n)^{4}.$
@@
To apply the bound in Lemma 4.1, we need to estimate the covering number $\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})$. The classical result of [22, Theorem 6] showed that the covering number can be bounded by pseudo-dimension:
$\log\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\lesssim\,{\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})\log(B_{n}/\epsilon),$ (4.3)
where ${\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})$ is the pseudo-dimension of the function class $\mathcal{T}_{B_{n}}\mathcal{F}_{n}$, see (2.2). For ReLU neural networks, [6] showed that
${\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})\lesssim N_{n}\log N_{n}.$
Consequently, we have
$\log\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\lesssim N_{n}\log(N_{n})\log(B_{n}/\epsilon).$
@@
$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{(\log n)^{2}N_{n}\log(N_{n})\log(nB_{n}^{2})}{n}+N_{n}^{-\frac{2\alpha}{d}}\lor M_{n}^{-\frac{4\alpha}{d+3-2\alpha}}.$
By choosing $N_{n}\asymp n^{d/(d+2\alpha)}$ and $M_{n}\gtrsim N_{n}^{(d+3-2\alpha)/(2d)}$, we get $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-2\alpha/(d+2\alpha)}(\log n)^{4}$. Similarly, by Lemma 4.1 and Lemma 2.3, if $\mathcal{H}=\mathcal{F}_{\sigma}(1)$ and $M_{n}\geq 1$, then
$\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim\frac{(\log n)^{2}N_{n}\log(N_{n})\log(nB_{n}^{2})}{n}+N_{n}^{-\frac{d+3}{d}}.$
We choose $N_{n}\asymp n^{d/(2d+3)}$, then $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-(d+3)/(2d+3)}(\log n)^{4}$. ∎
@@
The Gaussian noise assumption on the model (4.1) can be weaken for Lemma 4.1 and hence for Theorem 4.2. We refer the reader to [27, Appendix B, Lemma 18] for more details. Theorem 4.2 can be easily generalized to shallow ReLUk neural networks for $k\geq 1$ by using the same proof technique. For example, one can show that, if $h\in\mathcal{H}^{\alpha}$ with $\alpha<(d+2k+1)/2$, then we can choose $\mathcal{F}_{n}=\mathcal{F}_{\sigma_{k}}(N_{n},M_{n})$ with $N_{n}\asymp n^{\frac{d}{d+2\alpha}}$ and $M_{n}\gtrsim n^{\frac{d+2k+1-2\alpha}{2d+4\alpha}}$, such that $\mathbb{E}_{\mathcal{D}_{n}}\|\mathcal{T}_{B_{n}}f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}(\log n)^{4}$.
Theorem 4.2 shows that least square minimization using shallow ReLU neural networks can achieve the optimal rate $n^{-\frac{2\alpha}{d+2\alpha}}$ for learning functions in $\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$. For the function class $\mathcal{F}_{\sigma}(1)$, the rate $n^{-\frac{d+3}{2d+3}}$ is also minimax optimal as proven by [46, Lemma 25] (they studied a slightly different function class, but their result also holds for $\mathcal{F}_{\sigma}(1)$). Specifically, [57, Theorem 4 and Theorem 8] give a sharp estimate for the metric entropy
@@
Combining this estimate with the classical result of Yang and Barron (see [61, Proposition 1] and [60, Chapter 15]), we get
@@
where the infimum taken is over all estimators based on the samples $\mathcal{D}_{n}$, which are generated from the model (4.1).